how long will it take a principal to earn half of itself if invested at 12 percent simple interest?
Answer by fractalier(6550)

Normally we use the formula
I = prt
Since we want our interest to be equal to half our principal, we write
(1/2)p = prt
and plugging in we get
(1/2)p = p(.12)t
(1/2) = .12t
t = .5/.12 = 4 1/6 years